find the equation of the line parallel to the graph of 4x+2y=8 and containing (-1,5)
Answer by Abbey(339)

parallel lines have the same slope
put the equation into y-intercept form by subtracting 4x from both sides and then dividing by 2:
y=mx+b
y=-2x+4
so the parallel line slope will be -2
then plug the point(-1,5) into y=mx+b:
5=-2(-1)+b
5=2+b
3=b
so the equation that is parallel is:
y=-2x+3